# Wave propagation in an elastic lattice with non-reciprocal stiffness and engineered damping

### v1

@@ Abstract
Nonreciprocal wave propagation allows for directional energy transport. In this work, we systematically investigate wave dynamics in an elastic lattice that combines nonreciprocal stiffness with viscous damping. After establishing how conventional damping counteracts the system’s gain, we introduce a non-dissipative form of nonreciprocal damping in the form of gyroscopic damping. We find that the coexistence of nonreciprocal stiffness and nonreciprocal damping results in a decoupled control mechanism. The nonreciprocal stiffness is shown to govern the temporal amplification rate, while the nonreciprocal damper independently tunes the wave’s group velocity and oscillation frequency. This decoupling gives rise to phenomena such as the enhancement of net amplification for slower-propagating waves and boundary-induced wave mixing. These findings provide a theoretical framework for designing active metamaterials with more versatile control over their wave propagation characteristics.
@@ I Introduction
The paper is organized as follows. Section II reviews the model with nonreciprocal stiffness and analyzes its interaction with conventional damping. In Sec. III, we introduce a specialized, non-dissipative, nonreciprocal damper and examine its isolated effect on wave propagation. The full system combining both types of nonreciprocity is investigated in Sec. IV, where we detail the decoupled control mechanism and its consequences, such as boundary-induced wave mixing. Finally, Sec. V summarizes our findings and concludes the paper.
@@ II.1 No damper
[…]
@@
These expressions reveal that the leading-order correction appears exclusively in the imaginary part of the complex frequency, $\text{Im}[\omega(q)]$. Consequently, this term has a dominant effect on the temporal amplification or decay of the wave, while the corresponding correction to the real part of the frequency is of a higher order. It is noteworthy, however, that this small correction results in an increase in the oscillation frequency, a behavior opposite to that reported in some recently studied systems Rosa and Ruzzene (2020). Furthermore, for a fixed value of $\alpha$, the sign of $\text{Im}[\omega(q)]$ is determined by the sign of the wave number $q$.
@@
[…]
@@
$u_{n}(0)=\exp\left[-\frac{(n-n_{0})^{2}}{2\sigma^{2}}\right]\cos(q^{*}n),$
where $n_{0}=100$ is the center of the excitation, and $\sigma=12$ is the envelope width.
@@
Figure 2(a) shows the space-time evolution of the normalized displacement field, $u_{n}(t)/\|u(t)\|_{\infty}$, where $\|u(t)\|_{\infty}=\max_{n}|u_{n}(t)|$. An initial wave packet splits into two components traveling in opposite directions with the same group velocity. However, only the rightward-propagating component is amplified, a clear signature of the system’s non-reciprocal behavior. The group velocity measured from the slope of the propagating wavefront is $v_{g}\approx 7.184$, which is in close agreement with the slope calculated from the dispersion relation shown in Fig. 1(b).
To quantify the one-way amplification, in Fig. 2(b), we show the temporal evolution of the logarithm of the instantaneous global maximum of the lattice displacement field, $\log\left(\|\bm{u}(t)\|_{\infty}\right)$. The envelope of this logarithmic signal exhibits a linear trend, indicating exponential growth in the underlying displacement. A linear fit to the data yields a slope of approximately 0.71, representing the growth rate. This value closely matches the imaginary part of the complex dispersion relation, $\text{Im}[\omega(q^{*})]\approx 0.70$, shown in Fig. 1(c).
@@
Figure 2(c) shows the velocity time series for particles at positions $n=130$ and $n=160$, which are downstream from the excitation point. Both signals exhibit clear temporal amplification. The corresponding frequency spectra, presented in Fig. 2(d), reveal a dominant peak at $\omega\approx 24.38$. This value is consistent with the oscillation frequency, $\text{Re}[\omega(q^{*})]$, from Fig. 1(b), where $q^{*}$ is the wavenumber corresponding to the maximum growth rate.
@@
In contrast, the velocity signals from upstream positions ($n=49$ and $n=3$), plotted in Fig. 2(e), show significant decay over time. The associated frequency spectra in Fig. 2(f) contain the same dominant frequency component, but its amplitude is substantially reduced. This confirms that while the initial disturbance propagates in both directions, non-reciprocity leads to strong wave amplification in the forward (downstream) direction and attenuation in the backward (upstream) direction.
@@ II.2 Onsite Damper
$m\ddot{u}_{n}+c_{g}\dot{u}_{n}+(2k+k_{g})u_{n}\\ -k(1+\alpha)u_{n-1}-k(1-\alpha)u_{n+1}=0.$ (7)
@@
$\omega^{2}+i\frac{c_{g}}{m}\omega-\frac{1}{m}\Big{[}k_{g}+2k(1-\cos q)+\\ 2ik\alpha\sin q\Big{]}=0.$ (8)
@@
The dispersion relation for this system is plotted in Figs. 3(b) and 3(c). For small non-reciprocity ($\alpha\sim\mathcal{O}(\epsilon)$) and damping ($c_{g}\sim\mathcal{O}(\epsilon)$), it can be approximated by
@@
$-\frac{c_{g}^{2}}{8m^{2}\sqrt{A_{0}(q)}}+\mathcal{O}(\epsilon^{3}),$ (12)
@@
where $A_{0}(q)=\frac{1}{m}\left[k_{g}+2k(1-\cos q)\right]$. In contrast to the case with no damper, the imaginary part of the frequency is uniformly shifted downward by $-c_{g}/(2m)$, as seen in Fig. 3(c). Furthermore, the higher-order correction to the real part is now negative, causing a decrease in the oscillation frequency, which is opposite to the behavior without a damper. While $\text{Re}[\omega(q)]$ remains a symmetric function of $q$, the constant negative shift renders $\text{Im}[\omega(q)]$ asymmetric.
The parametric plots in Figs. 3(d) and 3(e) illustrate the interplay between non-reciprocity and damping. While the sign of group velocity is largely unchanged, the boundaries in the growth/decay map are significantly altered, leading to a reduction in the parameter space for amplification. A notable feature is the existence of traveling waves ($\text{Im}[\omega(q)]=0$) at wavenumbers away from $q=0$ and $q=\pm\pi$. By setting Eq. (13) to zero in the long-wavelength limit ($q\to 0$), we can find the condition for these waves:
@@
Figure 4: Numerical demonstration of directional wave amplification in a 1D lattice with nonreciprocal stiffness and onsite damping. (a) Space-time evolution of the normalized displacement field, showing amplified wave propagation in the forward direction with a measured group velocity $v_{g}\approx 7.184$. (b) Logarithmic plot of the instantaneous global maximum, where a linear fit yields a slope of 0.61, indicating a reduced temporal growth rate due to the presence of onsite damping. (c–d) Velocity responses and corresponding spectral amplitudes at downstream sites ($n=130,160$) confirm energy amplification and dominant excitation near $\omega\approx 24.38$. (e–f) Upstream sites ($n=49,3$) show significant attenuation in both time and frequency domains, affirming unidirectional energy transport.
@@
Figure 4(a) shows unidirectional wave propagation with a forward group velocity of $v_{g,f}\approx 7.18$, indicating that a small onsite damping does not significantly alter the wave speed. From Fig. 4(b), we measure a growth rate of 0.610.610.61, which confirms that the amplification persists, albeit at a reduced rate. The growth of velocity amplitudes at a downstream location, depicted in Figs. 4(c) and 4(d), further confirms this unidirectional amplification. This effect is weaker than in the undamped case; however, the spectral peak at $\omega\approx 24.38$ remains largely unaffected. Conversely, the backward-propagating signal shown in Figs. 4(e) and 4(f) is more strongly attenuated than its undamped counterpart, while its dominant frequency is unchanged. Overall, these findings demonstrate that the nonreciprocal amplification mechanism survives in the presence of a small onsite damping.
@@ II.3 Intersite Damper
$m\ddot{u}_{n}+c(2\dot{u}_{n}-\dot{u}_{n-1}-\dot{u}_{n+1})+(2k+k_{g})u_{n}\\ -k(1+\alpha)u_{n-1}-k(1-\alpha)u_{n+1}=0.$ (15)
@@
$\omega^{2}+i\frac{2c(1-\cos q)}{m}\omega\\ -\frac{1}{m}\left[k_{g}+2k(1-\cos q)+2ik\alpha\sin q\right]=0.$ (16)
@@
The dispersion relation for this system is plotted in Figs. 5(b) and 5(c). For small non-reciprocity ($\alpha\sim\mathcal{O}(\epsilon)$) and damping ($c\sim\mathcal{O}(\epsilon)$), we approximate the dispersion relation as
@@
$-\frac{c^{2}(1-\cos q)^{2}}{2m^{2}\sqrt{A_{0}(q)}}+\mathcal{O}(\epsilon^{3}),$ (20)
$\text{Im}[\omega(q)]$ $=\frac{k\alpha\sin q}{m\sqrt{A_{0}(q)}}-\frac{c(1-\cos q)}{m}+\mathcal{O}(\epsilon^{3}),$ (21)
where $A_{0}(q)=\frac{1}{m}\left[k_{g}+2k(1-\cos q)\right]$. In contrast to the uniform frequency shift from onsite damping, the imaginary part of the frequency in the intersite case is shifted downward by a wavenumber-dependent term, $-c(1-\cos q)/m$. As shown in Fig. 5(c), this damping effect is zero at $q=0$ and maximal at the edge of the Brillouin zone ($q=\pm\pi$). The higher-order correction to $\text{Re}[\omega(q)]$ due to damping is again negative, similar to the onsite case. Nevertheless, $\text{Re}[\omega(q)]$ still remains a symmetric function of $q$. However, the $q$-dependent shift results in a distinct asymmetric profile for $\text{Im}[\omega(q)]$.
We show the parametric plots in Figs. 5(d) and 5(e). While the group velocity characteristics remain largely unchanged, the boundaries separating gain and decay are qualitatively different from the onsite case. Notably, attenuation is now more pronounced near $q=\pm\pi$. Consequently, the condition for neutrally stable traveling waves ($\text{Im}[\omega]=0$) is reshaped. By setting the imaginary part of the frequency from our approximation to zero in the long-wavelength limit ($q\to 0$), we find this condition to be
@@
Figure 6: Numerical demonstration of directional wave amplification in a 1D lattice with nonreciprocal stiffness and intersite damping. (a) Space-time colormap of the normalized displacement field, showing rightward amplification of the wave packet. The measured group velocity $v_{g}\approx 7.249$ agrees with the theoretical prediction. (b) Logarithmic plot of the instantaneous global maximum of the displacement field, with a linear fit showing a slope of 0.45, in close agreement with the analytical growth rate. (c–d) Velocity profiles and corresponding spectra at downstream sites ($n=130,160$) confirm energy amplification at the dominant frequency $\omega\approx 24.38$. (e–f) Upstream responses at ($n=49,3$) show strong attenuation in both time and frequency domains.
We next examine the effect of intersite damping on the wave dynamics of the finite lattice. The results, presented in Fig. 6, show that unidirectional amplification of a wave packet is preserved, similar to the case with onsite damping. While the group velocity is comparable, the amplification rate differs significantly. This is consistent with the asymmetry observed in the imaginary part of the frequency, $\text{Im}(\omega)$, as shown in the dispersion diagram [Fig. 5(c)]. These findings demonstrate that the nonreciprocal amplification mechanism survives in the presence of a small intersite damping, although the growth rate is altered differently than for onsite damping.
To summarize, we have investigated the role of damping in a lattice with nonreciprocal stiffness by considering three representative cases: no damping, onsite damping, and intersite damping. The incorporation of onsite damping reduces the amplification rate uniformly across the wavenumber spectrum while preserving the directionality and coherence of the forward wave. In contrast, intersite damping introduces a wavenumber-dependent modification of the growth rates.

### v2

@@ Abstract
Nonreciprocal wave propagation allows for directional energy transport. In this work, we systematically investigate wave dynamics in an elastic lattice that combines nonreciprocal stiffness with viscous damping. After establishing how conventional damping counteracts the system’s gain, we introduce a non-dissipative form of nonreciprocal damping in the form of gyroscopic damping. We find that the coexistence of nonreciprocal stiffness and nonreciprocal damping results in a decoupled control mechanism. The nonreciprocal stiffness is shown to govern the temporal amplification rate, while the nonreciprocal damper independently tunes the wave’s group velocity and oscillation frequency. This decoupling gives rise to phenomena such as the enhancement of net amplification for slower-propagating waves, and also boundary-induced wave interference arising from divergent and convergent reflected wave trajectories with varying growth rates. These findings provide a theoretical framework for designing active metamaterials with more versatile control over their wave propagation characteristics.
@@ I Introduction
The paper is organized as follows. Section II reviews the model with nonreciprocal stiffness and analyzes its interaction with conventional damping. In Sec. III, we introduce a specialized, non-dissipative, nonreciprocal damper and examine its isolated effect on wave propagation. The full system combining both types of nonreciprocity is investigated in Sec. IV, where we detail the decoupled control mechanism and its consequence on tailoring the net amplification. We further demonstrate boundary-induced wave interference, showing how the interplay between stiffness and damping asymmetries can be engineered to combine frequency components. Finally, Sec. V summarizes our findings and concludes the paper.
@@ II.1 No damper
Figures 1(b) and 1(c) show the real and imaginary parts of the dispersion relation for $\alpha=0.1$, alongside the reciprocal case ($\alpha=0$) for comparison (we choose $k=180$ and $k_{g}=120$ as representative values for all simulations, as they clearly illustrate the phenomena and also yield operating frequencies on the same order of magnitude as recent experiments Brandenbourger et al. (2019)). The real part, $\text{Re}[\omega(q)]$, exhibits negligible variation, whereas the imaginary part, $\text{Im}[\omega(q)]$, reveals significant nonreciprocal behavior: attenuation for $q<0$ and amplification for $q>0$. These behaviors become clearer in the small-$\alpha$ limit, where the dispersion relation admits the following approximations:
@@
These expressions reveal that the leading-order correction appears exclusively in the imaginary part of the complex frequency, $\text{Im}[\omega(q)]$. Consequently, this term has a dominant effect on the temporal amplification or decay of the wave, while the corresponding correction to the real part of the frequency is of a higher order. It is noteworthy that this formulation results in an increase in the oscillation frequency for both positive and negative values of $\alpha$. This behavior contrasts with the findings of Rosa and Ruzzene Rosa and Ruzzene (2020), where a different stiffness definition ($k_{f}=k(1-\alpha)$ and $k_{b}=k$) was employed. Their definition leads to a decrease in frequency for positive $\alpha$ and an increase for negative $\alpha$.
@@
Figure 2: Finite-chain simulations of wave propagation and directional amplification in a lattice with nonreciprocal stiffness. (a) Space–time evolution of the space-normalized displacement field, $u_{n}(t)/\|u_{n}(t)\|_{\infty,n}$, in a 200-particle lattice with fixed boundaries showing selective amplification in the forward direction. The wavefront travels with group velocity $v_{g}\approx 7.184$. (b) Logarithmic plot of the instantaneous global maximum of the lattice displacement field, with a linear fit yielding a slope of 0.71. (c–d) Velocity responses and corresponding spectral amplitudes at downstream sites ($n=130,160$) highlight energy amplification and a spectral peak near $\omega\approx 24.38\,\text{rad/s}$. (e–f) Upstream sites ($n=70,40$) exhibit strong attenuation in both time and frequency domains, confirming unidirectional energy transport enabled by non-reciprocal stiffness.
@@
$u_{n}(0)=\exp\left[-\frac{(n-n_{0})^{2}}{2\sigma_{n}^{2}}\right]{{\cos(q^{*}(n-n_{0}))}},$
where $n_{0}=100$ is the center of the excitation, and ${\sigma_{n}}=12$ is the envelope width.
@@
Figure 2(a) shows the space-time evolution of the normalized displacement field, $u_{n}(t)/\|u_{n}(t)\|_{\infty,n}$, where $\|u_{n}(t)\|_{\infty,n}=\max_{n}|u_{n}(t)|$ denotes the spatial maximum of the displacement magnitude at time $t$. An initial wave packet splits into two components traveling in opposite directions with the same group velocity. However, only the rightward-propagating component is amplified, a clear signature of the system’s non-reciprocal behavior. The group velocity measured from the slope of the forward propagating wavefront is $v_{g}\approx 7.184$, which is in close agreement with the slope calculated from the dispersion relation shown in Fig. 1(b).
[…]
@@
Figure 2(c) shows the velocity time series for particles at positions $n=130$ and $n=160$, which are downstream from the excitation point. Both signals exhibit clear temporal amplification. The corresponding frequency spectra, presented in Fig. 2(d), reveal a spectral peak at $\omega\approx 24.38$. This value is consistent with the oscillation frequency, $\text{Re}[\omega(q^{*})]$, from Fig. 1(b), which corresponds to the central frequency of our narrow bandwidth Gaussian packet.
@@
In contrast, the velocity signals from upstream positions ($n=70$ and $n=40$), plotted in Fig. 2(e), show significant decay over time. The associated frequency spectra in Fig. 2(f) contain the same spectral peak, but its amplitude is substantially reduced. This confirms that while the initial disturbance propagates in both directions, non-reciprocity leads to strong wave amplification in the forward (downstream) direction and attenuation in the backward (upstream) direction.
Figure 3: Distinct trajectories of spatial and temporal peaks in a 1D finite lattice with nonreciprocal stiffness. (a) Analytical trajectories of the Gaussian envelope maxima, comparing the instantaneous spatial peaks $n_{\mathrm{peak}}(t)$ (dots and circles) and the temporal peaks $t_{\mathrm{peak}}(n)$ (solid lines). While the forward and backward $t_{\mathrm{peak}}(n)$ branches share the same slope (indicating equal effective group velocity), they exhibit distinct temporal intercepts at the source $n=n_{0}$, as highlighted in the inset. (b) Spatiotemporal evolution of the displacement field normalized by the local temporal maximum, $u_{n}(t)/\|u_{n}(t)\|_{\infty,t}$. The wavefronts propagate with equal speeds but project to unequal origin times, visually confirming the positive and negative intercepts predicted in (a).
It is interesting to note in Figs. 2(c) and 2(e) that the peak of the Gaussian time history occurs at different time instances, even though the observation points are located spatially symmetrically (e.g., equidistant upstream and downstream) from the initial excitation. This phenomenon is a specific feature of non-Hermitian systems governed by a complex dispersion relation.
For a narrow-bandwidth Gaussian wave packet with spatial variance $\sigma_{n}$, the trajectory of the spatial peak, $n_{\text{peak}}(t)$, is determined solely by the real part of the group velocity (see Appendix A for derivation):
$n_{\text{peak}}(t)=n_{0}+v_{g}\,t,$ (7)
where $v_{g}=\mathrm{Re}[\omega^{\prime}(q^{*})]$ denotes the standard group velocity evaluated at the carrier wavenumber $q^{*}$.
However, the evolution of the temporal peak, $t_{\text{peak}}(n)$—which corresponds to the maximum signal recorded by a stationary observer—depends on both the real and imaginary components of the dispersion. As derived in Appendix A, this relationship is given by:
$t_{\text{peak}}(n)=\frac{(n-n_{0})}{v_{g}^{\text{eff}}}+t_{\text{int}}\neq\frac{(n-n_{0})}{v_{g}},$ (8)
where the effective velocity is $v_{g}^{\text{eff}}=\mathrm{Re}[\omega^{\prime}]-(\mathrm{Im}[\omega^{\prime}])^{2}/\mathrm{Re}[\omega^{\prime}]$, and the intercept term is defined as:
$t_{\text{int}}=\frac{\sigma_{n}^{2}\,\mathrm{Im}[\omega(q^{*})]}{\big(\mathrm{Re}[\omega^{\prime}(q^{*})]\big)^{2}-\big(\mathrm{Im}[\omega^{\prime}(q^{*})]\big)^{2}}.$ (9)
Consequently, in a system with asymmetric complex dispersion, the forward and backward propagating packets may share the same group velocity $v_{g}^{\text{eff}}$ but possess different growth rates $\mathrm{Im}[\omega(q^{*})]$. For instance, as illustrated in Fig. 1(c), the forward-propagating wave (positive $q^{*}$) exhibits a positive growth rate ($\mathrm{Im}[\omega]>0$), which dictates a positive temporal intercept $t_{\text{int}}>0$. Conversely, the backward-propagating wave (negative $q^{*}$) is subject to attenuation ($\mathrm{Im}[\omega]<0$), leading to a negative intercept $t_{\text{int}}<0$. This difference, as visualized in Fig. 3(a), causes the temporal peaks at spatially symmetric locations to occur at different times. This behavior is further verified in Fig. 3(b), which plots the spatiotemporal map with time histories at each site normalized by their respective temporal peak displacement.
To avoid ambiguity between amplification and arrival times, the remaining sections of this work will focus on spatiotemporal maps showing the spatially normalized wave packet evolution. This approach isolates the transport characteristics, allowing us to relate the wave dynamics directly to the real group velocity $v_{g}$.
@@ II.2 Onsite Damper
$m\ddot{u}_{n}+c_{g}\dot{u}_{n}+(2k+k_{g})u_{n}-k(1+\alpha)u_{n-1}-k(1-\alpha)u_{n+1}=0.$ (10)
@@
$\omega^{2}+i\frac{c_{g}}{m}\omega-\frac{1}{m}\Big[k_{g}+2k(1-\cos q)+2ik\alpha\sin q\Big]=0.$ (11)
@@
The dispersion relation for this system is plotted in Figs. 4(b) and 4(c). For small non-reciprocity ($\alpha\sim\mathcal{O}(\epsilon)$) and damping ($c_{g}\sim\mathcal{O}(\epsilon)$), it can be approximated by
@@
$\quad-\frac{c_{g}^{2}}{8m^{2}\sqrt{A_{0}(q)}}+\mathcal{O}(\epsilon^{3}),$ (15)
@@
where $A_{0}(q)=\frac{1}{m}\left[k_{g}+2k(1-\cos q)\right]$. In contrast to the case with no damper, the imaginary part of the frequency is uniformly reduced by $c_{g}/(2m)$ without any dependency on $q$, as seen in Fig. 4(c). Furthermore, the higher-order correction to the real part is now negative, causing a decrease in the oscillation frequency, which is opposite to the behavior without a damper. While $\text{Re}[\omega(q)]$ remains a symmetric function of $q$, the constant negative shift renders $\text{Im}[\omega(q)]$ asymmetric.
The parametric plots in Figs. 4(d) and 4(e) illustrate the interplay between non-reciprocity and damping. While the sign of group velocity is largely unchanged, the boundaries in the growth/decay map are significantly altered, leading to a reduction in the parameter space for amplification. A notable feature is the existence of traveling waves ($\text{Im}[\omega(q)]=0$) at wavenumbers away from $q=0$ and $q=\pm\pi$. By setting Eq. (16) to zero in the long-wavelength limit ($q\to 0$), we can find the condition for these waves:
@@
Figure 5: Finite-chain simulations of wave propagation and directional amplification in a lattice with nonreciprocal stiffness and onsite damping. (a) Space–time evolution of the space-normalized displacement field, showing amplification in the forward-propagating wave packet with measured group velocity $v_{g}\approx 7.184$. (b) Logarithmic plot of the instantaneous global maximum, where a linear fit yields a slope of 0.61, indicating a reduced temporal growth rate due to the presence of onsite damping. (c–d) Velocity responses and corresponding spectral amplitudes at downstream sites ($n=130,160$) confirm energy amplification and peak excitation near $\omega\approx 24.38$. (e–f) Upstream sites ($n=70,40$) show significant attenuation in both time and frequency domains, affirming unidirectional energy transport.
@@
Figure 5(a) shows unidirectional wave propagation with a forward group velocity of ${v_{g}}\approx 7.18$, indicating that a small onsite damping does not significantly alter the wave speed. From Fig. 5(b), we measure a growth rate of $0.61$, which confirms that the amplification persists, albeit at a reduced rate. The growth of velocity amplitudes at a downstream location, depicted in Figs. 5(c) and 5(d), further confirms this unidirectional amplification. This effect is weaker than in the undamped case; however, the spectral peak at $\omega\approx 24.38$ remains largely unaffected. Conversely, the backward-propagating signal shown in Figs. 5(e) and 5(f) is more strongly attenuated than its undamped counterpart, while its peak frequency is unchanged.
Overall, these findings demonstrate that the nonreciprocal amplification mechanism and its kinematic signatures survive in the presence of a small onsite damping.
II.3 Intersite Damper
@@
$\begin{split}m\ddot{u}_{n}&+c(2\dot{u}_{n}-\dot{u}_{n-1}-\dot{u}_{n+1})\\ &+(2k+k_{g})u_{n}-k(1+\alpha)u_{n-1}-k(1-\alpha)u_{n+1}=0.\end{split}$ (18)
@@
$\begin{split}\omega^{2}&+i\frac{2c(1-\cos q)}{m}\omega\\ &-\frac{1}{m}\left[k_{g}+2k(1-\cos q)+2ik\alpha\sin q\right]=0.\end{split}$ (19)
@@
The dispersion relation for this system is plotted in Figs. 6(b) and 6(c). For small non-reciprocity ($\alpha\sim\mathcal{O}(\epsilon)$) and damping ($c\sim\mathcal{O}(\epsilon)$), we approximate the dispersion relation as
@@
$\quad-\frac{c^{2}(1-\cos q)^{2}}{2m^{2}\sqrt{A_{0}(q)}}+\mathcal{O}(\epsilon^{3}),$ (23)
$\text{Im}[\omega(q)]$ $=\frac{k\alpha\sin q}{m\sqrt{A_{0}(q)}}-\frac{c(1-\cos q)}{m}+\mathcal{O}(\epsilon^{3}).$ (24)
where $A_{0}(q)=\frac{1}{m}\left[k_{g}+2k(1-\cos q)\right]$. In contrast to the uniform decrease in frequency for the onsite case in Sec. IIB, the imaginary part of the frequency in the intersite case is reduced by a wavenumber-dependent term, $c(1-\cos q)/m$. As shown in Fig. 6(c), this damping effect is zero at $q=0$ and maximal at the edge of the Brillouin zone ($q=\pm\pi$). The higher-order correction to $\text{Re}[\omega(q)]$ due to damping is again negative, similar to the onsite case. Nevertheless, $\text{Re}[\omega(q)]$ still remains a symmetric function of $q$. However, the $q$-dependent shift results in a distinct asymmetric profile for $\text{Im}[\omega(q)]$.
We show the parametric plots in Figs. 6(d) and 6(e). While the group velocity characteristics remain largely unchanged, the boundaries separating gain and decay are qualitatively different from the onsite case. Notably, attenuation is now more pronounced near $q=\pm\pi$. Consequently, the condition for neutrally stable traveling waves ($\text{Im}[\omega]=0$) is reshaped. By setting the imaginary part of the frequency from our approximation to zero in the long-wavelength limit ($q\to 0$), we find this condition to be
@@
Figure 7: Finite-chain simulations of wave propagation and directional amplification in a lattice with nonreciprocal stiffness and intersite damping. (a) Space–time evolution of the space-normalized displacement field, showing rightward amplification of the wave packet. The measured group velocity $v_{g}\approx 7.249$ agrees with the theoretical prediction. (b) Logarithmic plot of the instantaneous global maximum of the displacement field, with a linear fit showing a slope of 0.45, in close agreement with the analytical growth rate. (c–d) Velocity profiles and corresponding spectra at downstream sites ($n=130,160$) confirm energy amplification at the peak frequency $\omega\approx 24.38$. (e–f) Upstream responses at ($n=70,40$) show strong attenuation in both time and frequency domains.
We next examine the effect of intersite damping on the wave dynamics of the finite lattice. The results, presented in Fig. 7, show that unidirectional amplification of a wave packet is preserved, similar to the case with onsite damping. While the group velocity is comparable, the amplification rate differs significantly. This is consistent with the asymmetry observed in the imaginary part of the frequency, $\text{Im}(\omega)$, as shown in the dispersion diagram [Fig. 6(c)]. These findings demonstrate that the nonreciprocal amplification mechanism survives in the presence of a small intersite damping, although the growth rate altered differently than for onsite damping.
To summarize, we have investigated the role of damping in a lattice with nonreciprocal stiffness by considering three representative cases: no damping, onsite damping, and intersite damping. The effect of onsite damping is independent of the wavenumber $q$. As seen from Eq. (16), it contributes a constant term of $c_{g}/2m$ to $\text{Im}(\omega)$, leading to uniform attenuation of all modes. In contrast, intersite damping introduces a wavenumber-dependent term of the form $c(1-\cos q)/m$ (see Eq. (24)). This variation produces selective attenuation across the Brillouin zone, modifying the growth and decay rates non-uniformly.